a mother is twice as old as her daughter now. sixteen years ago,she was four times as old. how old is each now?

Answer by Maths68(1474)

Let
Present age of mother = m
Present age of daughter = d
Mother is twice as old as her daughter now.
m=2d.......(1)
sixteen years ago,she was four times as old
m-16=4(d-16)
m-16=4d-64
m=4d-64+16
m=4d-48.........(2)
subtitute the value of m from (1) to (2)
2d=4d-48
2d-4d=-48
-2d=-48
-2d/-2=-48/-2
d=24
Put the value of d in (1)
m=2d
m=2*24
m=48
resent age of mother = m = 48 years old
Present age of daughter = d = 24 years old